OpenStudy (anonymous):

Help simplifying:

OpenStudy (anonymous):

\[\frac{ 2x^3+4x^2+10x+4 }{ x^3+3x^2+5x+3 }\]

OpenStudy (goldphenoix):

 Does it make sense now? (Btw, this isn't in simpliest form. You know what to do next?

OpenStudy (goldphenoix):

\[\large \large \large \large \large \frac{ 2x^3 }{ x^3 } + \frac{ 4x^2 }{ 3x^2 } + \frac{ 10x }{ 5x } + \frac{ 4 }{ 3 }\] ^Just to make it neater. >_> So you know what to do next?

OpenStudy (anonymous):

2+1+2+4/3?

OpenStudy (goldphenoix):

^How did you get that?

OpenStudy (anonymous):

I don't think I'm getting it

OpenStudy (goldphenoix):

Do you understand how I split it into 4 monomials?

OpenStudy (anonymous):

Yes. What's 2x^3/x^3? Doesn't the x^3 cancel out?

OpenStudy (goldphenoix):

Well. x^3 divided by x^3 cancel out. 2 is still there. So it would be 2. Now what's 4x^2 divided by 3x^2?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

I mean x^2

OpenStudy (goldphenoix):

Remember. We're dividing. x^2 divided by x^2 cancel out. We're left with: \[\large \large \frac{ 4 }{ 3 } OR 1.33\]

OpenStudy (goldphenoix):

What would the next monomial simplify to?

OpenStudy (anonymous):

Okay. 2

OpenStudy (goldphenoix):

Correct!!! What's the next one?

OpenStudy (anonymous):

Just 4/3

OpenStudy (anonymous):

So the answer is 2+4/3+2+4/3?

OpenStudy (goldphenoix):

No.

OpenStudy (goldphenoix):

You remember each simplify monomial gives you, right? You basically add up all of them. (Since they're all like terms) What would that give you?

OpenStudy (anonymous):

12/6 2

OpenStudy (goldphenoix):

It should give you: 2 + (4/3) + 2 + (4/3) What does that simplify to? Well 2 + 2 is 4. (4/3) + (4/3) = (8/3) or (2.66) So your answer is either 20/3 or 6.66 How did I get 20/3? Well, 4 can be written as (4/1) To add (4/1) to (8/3), you want to find the least common denominator. I can multiply (4/1) by 3 to get (12/3). (12/3) and (8/3) has a common denominator. The final step to add both fraction is to add the numerator, but keep the denominator. This will give you 20/3. I hope this help!

OpenStudy (anonymous):

Ah. Thank you so much!

OpenStudy (goldphenoix):

No problem. :) As long as you understand it, those problem would be quite easy for you. Just keep on practicing. :)